OpenStudy (cam1258):

HELP ASAP!!!!!!!!!! What are the values of a, b, and c in the quadratic equation 0 =1/2x² – 3x – 2? A)a = , b = 3, c = 2 B)a = , b = –3, c = –2 C)a = , b = 3, c = –2 D)a = , b = –3, c = 2

OpenStudy (shadowlegendx):

\[Ax + By + C = 0\]

jhonyy9 (jhonyy9):

please use parentheses this 1/2 is the coeficient of x^2 ?

OpenStudy (cam1258):

1/2(^2)

jhonyy9 (jhonyy9):

@ShadowLegendX i think you thought on Ax^2 +Bx +C = 0 yes ?

OpenStudy (shadowlegendx):

\[\frac{ 1 }{ 2 }x^2 - 3x - 2 = 0\] \[Ax^2 + Bx + C = 0\]

OpenStudy (shadowlegendx):

Yes

OpenStudy (cam1258):

im running out of time

OpenStudy (cam1258):

please help

OpenStudy (cam1258):

@jhonyy9

OpenStudy (cam1258):

@ShadowLegendX

jhonyy9 (jhonyy9):

1 ---x2−3x−2=0 2 Ax^2 +Bx+C =0

jhonyy9 (jhonyy9):

what is the coeficient of x^2 ? so this mean the A